OpenStudy (anonymous):

5. Essay; show all work. A simple dartboard has three areas… the main board has a radius of 8 inches, there is a circle with a radius of 5 inches, and the bullseye has a radius of 2 inches. What is the probability of a random dart landing inside the bullseye?

OpenStudy (anonymous):

2^2 pi/8^2 pi=4/64=1/16